# Sublinear bilipschitz equivalence and the quasiisometric classification of solvable Lie groups

### v1

@@ 3.A. Some preliminaries on distortion and Cornulier’s class
    $H$ is a $O(\log)$-Lipschitz retract of $G$, more precisely:
    1. (a)
      $\pi:G\to H$ is $O(\log)$-Lipschitz;
    2. (b)
      Let $X$ be a nonzero vector in a Cartan subalgebra of $\mathfrak{g}$. Then there exists $f:H\to G$ (depending on $X$) which is $O(\log)$-Lipschitz and such that
      1. (i)
        $\pi\circ f$ is $O(\log)$-close to the identity of $H$.
      2. (ii)
        $f\circ\pi(\exp(tX))=\exp(tX)$ for all $t$ in $\mathbf{R}$.
@@
Proof.
Parts (1) and (3) are stated by Cornulier [Cor11]. Parts (2a) and (2bi) express that $\pi\colon G\to H$ is a retract in the category of $O(\log)$-Lipschitz maps, which is also stated in [Cor19, Example 2.6] and the content of the proof can be found [Cor11, Theorem 4.4], where in the notation of [Cor11], $f$ is the map $\psi^{-1}_{\mid V}$ (before the statement of Theorem 4.4). See also the few lines before [Cor08, Lemma 5.2] where the map $f$ that we need is named $\psi$. To check part (2bii) we have to specify the construction of $f$; for this we refer to some parts of Cornulier’s proof in [Cor11]. In Cornulier’s construction, $\pi(X)$ is identified with an element $\xi$ in a subspace of $V$, the complement of the Lie algebra of $\mathfrak{h}\cap\operatorname{R}_{\exp}\mathfrak{g}$ in $\mathfrak{h}$ where $\mathfrak{h}$ can be taken to be any Cartan subalgebra of $\mathfrak{g}$, and then, $f\circ\pi(\exp(tX))=\exp_{G}(t\xi)$. Since in our assumption, $X$ lies in a Cartan subalgebra $\mathfrak{h}$ of $X$, we can take the Cartan subalgebra in the construction of $f$ to be $\mathfrak{h}$, and then, with this choice, take $\xi$ to be equal to $X$. In this way, $f\circ\pi(\exp(tX))=\exp(tX)$ for all $t\in\mathbf{R}$. ∎
The following notion will be useful for us, as it allows to use the work of Cornulier and Tessera for upper bounds on Dehn functions.
Definition 3.8 (After Cornulier and Tessera, [CT17, Definition 1.2]).
[…]
Proposition 3.9.
Let $G$ be a completely solvable group. Assume that $G$ splits as $U\rtimes A$, where $A$ is abelian, $U$ is nilpotent, and the action of $A$ on $U/[U,U]$ has no fixed point. Then
$U=[G,G]=\operatorname{R}_{\exp}G.$
Moreover, the action of $A$ on any non-trivial quotient of $U/[U,U]$ has no fixed point, hence $G$ is standard solvable in the sense of [CT17].
Proof.
With notation as before, consider
$\mathfrak{g}=\mathfrak{u}\rtimes\mathfrak{a},$
and denote $\rho:\mathfrak{a}\to\mathfrak{gl}(\mathfrak{u}/[\mathfrak{u},\mathfrak{u}])$. In order to prove that $\mathfrak{u}\subset[\mathfrak{g},\mathfrak{g}]$, it is enough to prove that $\rho(\mathfrak{a})(\mathfrak{u}/[\mathfrak{u},\mathfrak{u}])=\mathfrak{u}/[\mathfrak{u},\mathfrak{u}]$. Indeed, if for $u\in\mathfrak{u}$, we can write
$u=[a,u^{\prime}]+w$
[…]
We proved that $\mathfrak{u}\subseteq[\mathfrak{g},\mathfrak{g}]$. The converse containment follows from the fact that $\mathfrak{g}/\mathfrak{u}=\mathfrak{a}$ is abelian. Now, $C^{3}\mathfrak{g}=[\mathfrak{g},\mathfrak{g}]$ because of the following series of equalities
$[\mathfrak{g},\mathfrak{u}]=[\mathfrak{a}+\mathfrak{u},\mathfrak{u}]=[\mathfrak{a},\mathfrak{u}]+[\mathfrak{u},\mathfrak{u}]=\mathfrak{u}.$
and in view of the fact that $\operatorname{R}_{\exp}\mathfrak{g}$ is the limit of the central series, $\mathfrak{u}=\operatorname{R}_{\exp}{\mathfrak{g}}$. ∎
We can check that all the groups of class $(\mathcal{C}_{0})$ of dimension less than $5$, save for one, are standard solvable (See Table 4). The only exception is the group $G_{4,3}$, since $G_{4,3}/\operatorname{R}_{\exp}G_{4,3}$ is non-abelian. There are many non-standard solvable groups of dimension 5 and class $(\mathcal{C}_{0})$ see Section 5.C.2.
@@
Let’s check that $\mathbf{R}\times G_{3,3}$ is the group in the class $(\mathcal{C}_{1})$ associated to $G_{4,9}^{0}$ by [Cor11], and so there is a $O(\log)$-sublinear bilipschitz equivalence $\phi\colon G_{4,9}^{0}\to\mathbf{R}\times G_{3,3}$. The Lie algebra $\mathfrak{g}=\mathfrak{g}_{4,9}^{0}$ of $G_{4,9}^{0}$ is a semidirect product $\mathfrak{heis}\rtimes\delta$ where $\mathfrak{heis}$ has a basis $(e_{1},e_{2},e_{3})$ with $[e_{1},e_{2}]=e_{3}$ and $\delta e_{i}=e_{i}$ if $i=1,3$ or 00 if $i=2$. The derived subalgebra is $\mathfrak{u}=\langle e_{1},e_{3}\rangle$ and $[\mathfrak{g},\mathfrak{u}]=\mathfrak{u}$, so that $\mathfrak{u}$ is the Lie algebra of the exponential radical. The matrices of $\operatorname{ad}_{e_{3}}$ and $\operatorname{ad}_{e_{4}}$ in the basis $(e_{1},e_{2})$ are, respectively,
@@ 3.C. Proof of Corollary B and Theorem C
Some Heintze groups of diagonal type are parabolic subgroups of isometries of rank one symmetric spaces; they do not have the strong pointed sphere property, since the group of isometries of a rank one symmetric space acts transitively on its Gromov boundary. Among the Heintze groups with an abelian nilradical, the ones of the latter kind turn out to be the only exceptions:
@@
Let $m$, $n$, $m^{\prime}$ and $n^{\prime}$ be nonnegative integers. Let $G$, and $G^{\prime}$ be two real semisimple Lie groups with trivial center and no compact factors. Let $\{H_{i}\}_{1\leqslant i\leqslant m}$ and $\{H^{\prime}_{j}\}_{1\leqslant j\leqslant m^{\prime}}$ be families of diagonal Heintze groups satisfying the strong pointed sphere property; let $M_{i}$, resp. $M^{\prime}_{j}$ be the nilradical of $H_{i}$, resp. of $H^{\prime}_{j}$. Write $G=KAN$, and $G^{\prime}=K^{\prime}A^{\prime}N^{\prime}$. Form the following solvable groups $P$, $P^{\prime}$, $R\triangleleft P$ and $R^{\prime}\triangleleft P^{\prime}$
$P=\mathbf{R}^{n}\times AN\times\prod_{i=1}^{m}H_{i}$ $R=N\times\prod_{i=1}^{m}M_{i};$
$P^{\prime}=\mathbf{R}^{n^{\prime}}\times A^{\prime}N^{\prime}\times\prod_{j=1}^{m^{\prime}}H^{\prime}_{j}$ $R^{\prime}=N^{\prime}\times\prod_{j=1}^{m^{\prime}}M^{\prime}_{j}.$
@@
The strong pointed sphere property is known to hold for some Heintze groups with nonabelian nilradicals [Pal22, Remark 9]. So the assumptions of Theorem 3.17 are indeed weaker than those of Theorem C.
@@
where $G_{i}$ has $\mathbf{R}$-rank one for $1\leqslant i\leqslant p$ and $\mathbf{R}$-rank at least two for $p+1\leqslant i\leqslant p+q$, similarly for $G^{\prime}_{j}$. Let $\ell=p+m$ and $\ell^{\prime}=p^{\prime}+m^{\prime}$. For $i\in\{m+1,\ell\}$ define $M_{i}:=N_{i-m}$ and $H_{i}:=A_{i-m}N_{i-m}$, where $G_{k}=K_{k}A_{k}N_{k}$ for $1\leqslant k\leqslant p$; similarly, define $M^{\prime}_{j}$ and $H^{\prime}_{j}$ for $j$ in $\{m^{\prime}+1,\ell^{\prime}\}$. In this way, $\{H_{i}\}_{i=1}^{m}$, resp. $\{H_{j}\}_{j=1}^{m^{\prime}}$ are the Heintze groups that were originally present in the statement, and $\{H_{i}\}_{i=m+1}^{m+p}$, resp. $\{H_{j}\}_{j=m^{\prime}+1}^{m^{\prime}+p^{\prime}}$ are the Borel subgroups of the rank $1$ factors of $G$, resp. $G^{\prime}$ (which are also Heintze groups).
It follows from Theorem 3.7(1) that the groups $S$ and $P$ on the one hand, $S^{\prime}$ and $P^{\prime}$ on the other hand, are $O(\log)$-bilipschitz equivalent. By assumption, $S$ and $S^{\prime}$ are quasiisometric, especially they are $O(\log)$-bilipschitz equivalent. So $P$ and $P^{\prime}$ are $O(\log)$-bilipschitz equivalent.
Applying Theorem A to $P$ and $P^{\prime}$ we find that $n=n^{\prime}$, $q=q^{\prime}$, $\ell=\ell^{\prime}$, and there are two bijections $\sigma_{1}$ of $\{1,\ldots,\ell\}$ and $\sigma_{2}$ of $\{p+1,\ldots p+q\}$ so that
  1. (1)
    $G_{i}$ is $O(\log)$-bilipschitz equivalent to $G_{\sigma_{2}(i)}$ for all $p<i\leqslant p+q$;
  2. (2)
    $H_{i}$ is $O(\log)$-bilipschitz equivalent to $H_{\sigma_{1}(i)}$ for all $1\leqslant i\leqslant\ell$;
From (1), by Lemma 3.15 we deduce that the groups
@@
are isomorphic. We now claim that $\sigma_{1}(\{1,\ldots,p\})=\{1,\ldots,p^{\prime}\}$, especially $p=p^{\prime}$. If it was not the case, then some $H_{i}$, say $H_{i_{0}}$, would be $O(u)$-equivalent to a rank one symmetric space. But then the group of self-SBEs of $H_{i}$ would be transitive on its sphere at infinity. This cannot be, as the pointed sphere conjecture holds for $H_{i}$ by assumption. It now follows from (2) and [Pal20a] that the groups

### v2

@@ 3.A. Some preliminaries on distortion and Cornulier’s class
    $H$ is a $O(\log)$-Lipschitz retract of $G$, more precisely: $\pi:G\to H$ is $O(\log)$-Lipschitz.
@@
Let’s check that $\mathbf{R}\times G_{3,3}$ is the group in the class $(\mathcal{C}_{1})$ associated to $G_{4,9}^{0}$ by [Cor11], and so there is a $O(\log)$-sublinear bilipschitz equivalence $\phi\colon G_{4,9}^{0}\to\mathbf{R}\times G_{3,3}$. The Lie algebra $\mathfrak{g}=\mathfrak{g}_{4,9}^{0}$ of $G_{4,9}^{0}$ is a semidirect product $\mathfrak{heis}\rtimes\delta$ where $\mathfrak{heis}$ has a basis $(e_{1},e_{2},e_{3})$ with $e_{1},e_{2}]=e_{3}$ and $\delta e_{i}=e_{i}$ if $i=1,3$ or 0 if $i=2$. The derived subalgebra is $\mathfrak{u}=\langle e_{1},e_{3}\rangle$ and $\mathfrak{g},\mathfrak{u}]=\mathfrak{u}$, so that $\mathfrak{u}$ is the Lie algebra of the exponential radical. The matrices of $\operatorname{ad}_{e_{3}}$ and $\operatorname{ad}_{e_{4}}$ in the basis $(e_{1},e_{2})$ are, respectively,
@@ 3.C. Proof of Corollary B and Theorem C
Some examples of Heintze groups of diagonal type come from rank one symmetric spaces. More generally, we define Iwasawa subgroups as follows.
Definition 3.11.
Let $G$ be a simple Lie group with trivial center. Let $G=KAN$ be an Iwasawa decomposition of $G$. We call $AN$ an Iwasawa subgroup of $G$; it is unique up to conjugation in $G$ so that we may speak of “the Iwasawa subgroup of GG”. If $G$ has real rank one, its Iwasawa subgroup is a Heintze group of diagonal type, see [Pan89, §9.3].
@@
The Iwasawa subgroups of simple Lie groups of rank one do not have the strong pointed sphere property, since the group of isometries of a rank one symmetric space acts transitively on its Gromov boundary. Among the Heintze groups with an abelian nilradical, the ones of the latter kind turn out to be the only exceptions:
@@
Let $m$, $n$, $m^{\prime}$ and $n^{\prime}$ be nonnegative integers. Let $G$, and $G^{\prime}$ be two real semisimple Lie groups with trivial center and no compact factors. Let $\{H_{i}\}_{1\leqslant i\leqslant m}$ and $\{H^{\prime}_{j}\}_{1\leqslant j\leqslant m^{\prime}}$ be families of diagonal Heintze groups satisfying the strong pointed sphere property. Write $G=KAN$, and $G^{\prime}=K^{\prime}A^{\prime}N^{\prime}$. Form the following solvable groups $P$ and $P^{\prime}$ :
$P=\mathbf{R}^{n}\times AN\times\prod_{i=1}^{m}H_{i}\;\text{and}\ P^{\prime}=\mathbf{R}^{n^{\prime}}\times A^{\prime}N^{\prime}\times\prod_{j=1}^{m^{\prime}}H^{\prime}_{j}.$
@@
The strong pointed sphere property is known to hold for some Heintze groups with nonabelian nilradicals [Pal22, Remark 9]. So the assumptions of Theorem 3.16 are indeed weaker than those of Theorem C. (See also Remark 4.5 for concrete examples.)
@@
We now come to the proof of Theorem 3.16. Applying Theorem A to $P$ and $P^{\prime}$ we will get $O(\log)$-equivalences between the direct factors of $P$ and $P^{\prime}$. The proof consists in establishing which factors can be paired with one another. While it requires some notation, the reality is quite clear: simple factors of $G$ and $G^{\prime}$ must pair to one another, whilst preserving the $\mathbf{R}$-rank one vs. higher $\mathbf{R}$-rank distinction; the strong pointed sphere property further allows to ensure that the remaining factors, namely the Heintze groups of diagonal type with this property, are not paired with the $\mathbf{R}$-rank one factors which do not have it.
@@
where $G_{i}$ has $\mathbf{R}$-rank one for $1\leqslant i\leqslant p$ and $\mathbf{R}$-rank at least two for $p+1\leqslant i\leqslant p+q$, similarly for $G^{\prime}_{j}$. Let $\ell=p+m$ and $\ell^{\prime}=p^{\prime}+m^{\prime}$. For $i\in\{m+1,\ell\}$ define $H_{i}:=A_{i-m}N_{i-m}$, where $G_{k}=K_{k}A_{k}N_{k}$ for $1\leqslant k\leqslant p$; similarly, define $H^{\prime}_{j}$ for $j$ in $\{m^{\prime}+1,\ell^{\prime}\}$. In this way, $\{H_{i}\}_{i=1}^{m}$, resp. $\{H_{j}\}_{j=1}^{m^{\prime}}$ are the Heintze groups that were originally present in the statement, and $\{H_{i}\}_{i=m+1}^{m+p}$, resp. $\{H_{j}\}_{j=m^{\prime}+1}^{m^{\prime}+p^{\prime}}$ are the Iwasawa subgroups of the rank $1$ factors of $G$, resp. $G^{\prime}$ (which are also diagonal type Heintze groups). Especially, the groups $H_{i}$ for $1\leqslant i\leqslant m$ and $H^{\prime}_{j}$ for $1\leqslant j\leqslant m^{\prime}$ have coarse type I.
It follows from Theorem 3.7(1) that the groups $S$ and $P$ on the one hand, $S^{\prime}$ and $P^{\prime}$ on the other hand, are $O(\log)$-bilipschitz equivalent. By assumption, $S$ and $S^{\prime}$ are quasiisometric, especially they are $O(\log)$-bilipschitz equivalent. So $P$ and $P^{\prime}$ are $O(\log)$-bilipschitz equivalent. Note that the coarse type I factors of $P$, resp. $P^{\prime}$ are the groups $H_{i}$, resp. the groups $H^{\prime}_{j}$, while the coarse type II factors are of $P$, resp. $P^{\prime}$ are the groups $A_{k}N_{k}$ for $k\in\{p+1,\ldots,q\}$, resp. $A^{\prime}_{s}N^{\prime}_{s}$ for $s\in\{p^{\prime}+1,\ldots,p^{\prime}+q^{\prime}\}$, which are respectively quasiisometric to the simple Lie groups $G_{k}$ and $G^{\prime}_{s}$ of higher $\mathbf{R}$-rank.
Applying Theorem A to $P$ and $P^{\prime}$ we find that $n=n^{\prime}$, $q=q^{\prime}$, $\ell=\ell^{\prime}$, and there are two bijections $\sigma_{\mathrm{I}}$ of $\{1,\ldots,\ell\}$ and $\sigma_{\mathrm{II}}$ of $\{p+1,\ldots p+q\}$ so that
  1. (I)
    $H_{i}$ is $O(\log)$-bilipschitz equivalent to $H_{\sigma_{\mathrm{I}}(i)}$ for all $1\leqslant i\leqslant\ell$;
  2. (II)
    $G_{i}$ is $O(\log)$-bilipschitz equivalent to $G_{\sigma_{\mathrm{II}}(i)}$ for all $p<i\leqslant p+q$;
From (II) and Lemma 3.14 we deduce that the groups
@@
are isomorphic. We now claim that $\sigma_{\mathrm{I}}(\{1,\ldots,p\})=\{1,\ldots,p^{\prime}\}$, especially $p=p^{\prime}$. If it was not the case, then some $H_{i}$, say $H_{i_{0}}$, with $i_{0}>p$, would be $O(u)$-equivalent to a rank one symmetric space. But then the group $\operatorname{SBE}(H_{i_{0}})^{O(\log)}$ would be transitive on its Gromov boundary. This cannot be, as $H_{i_{0}}$ has the strong pointed sphere property. It now follows from (I) and [Pal20a] that the factors of